# Euclid: Early Release Observations – The star cluster systems of the Local Group dwarf galaxies IC 10 and NGC 6822††thanks: This paper is published on behalf of the Euclid Consortium

### v1

@@
Star clusters are valuable indicators of galaxy evolution, offering insights into the buildup of stellar populations across cosmic time. Understanding the intrinsic star cluster populations of dwarf galaxies is particularly important given these systems’ role in the hierarchical growth of larger systems. We use data from Euclid’s Early Release Observation programme to study star clusters in two star-forming dwarf irregular galaxies in the Local Group, NGC 6822 and IC 10 [ (1–4) ]. With Euclid, star clusters are resolved into individual stars across the main bodies and haloes of both galaxies. Through visual inspection of the images, we uncover 30 new star cluster candidates in NGC 6822 and 16 in IC 10, ranging from compact to diffuse extended clusters. We compile and re-evaluate previously identified literature candidates, resulting in final combined catalogues of 52 (NGC 6822) and 71 (IC 10) cluster candidates with confidence-based classifications. We present homogeneous photometry in , , , and , and in archival UBVRI data, alongside size measurements and properties derived from the SED-fitting code BAGPIPES. Through synthetic cluster injection, we conclude our sample is complete to for ages , and to for ages of . We find that IC 10 has more young clusters than NGC 6822, and its young clusters extend to higher masses, consistent with its starburst nature. We find several old massive () clusters in both dwarfs, including an exceptional cluster in NGC 6822’s outskirts with a mass of , nearly twice as massive as any other old cluster in either galaxy. In NGC 6822, we also identify a previously undetected, old, and extended cluster (). Using well-defined criteria, we identify 11 candidate GCs in NGC 6822 and eight in IC 10. Both galaxies have high specific frequencies () for their luminosities but remain consistent with the known GC scaling relationships in the low-luminosity regime.
@@
Star clusters are present in nearly all galaxies and can be observed across most of cosmic time (e.g., Forbes et al. 2018a; Adamo et al. 2024). They are likely the sites of most, if not all, star formation in the Universe, even if they do not always survive for long as coherent or bound groupings of stars. In the Local Volume, star clusters are observed with a remarkable range of properties, from massive and young clusters (e.g., Portegies Zwart et al. 2010) to ancient, faint, and extended ones (e.g., Huxor et al. 2005; Crnojević et al. 2016). The origin of this diversity is not fully understood but likely reflects the complex interplay between birth conditions, stellar feedback, and dynamical evolution (e.g., Krumholz et al. 2019; Lahén et al. 2025). Nearby galaxies offer key advantages for studying star clusters, as their proximity allows resolution into individual stars, facilitating both identification and characterisation (e.g., Huxor et al. 2014; Johnson et al. 2017; Jang et al. 2012). Furthermore, using resolved stars, star clusters can be observed to lower luminosities and surface densities than is possible in more distant systems, yielding insight into small-scale star-formation events and the bottom end of the cluster mass function.
@@
Comprehensive views of star cluster populations in nearby galaxies are still relatively rare. Historically, many extragalactic star cluster studies have been conducted with the Hubble Space Telescope (HST), which captures only a small field of view (FoV) in a single pointing. Studies often choose to focus on the bright inner regions of galaxies where recent star formation is concentrated and hence the number of young clusters is highest (e.g., Adamo et al. 2017). Even large HST imaging campaigns, such as the PHAT and PHATTER surveys of M31 and M33, capture star clusters only in the main disc (Johnson et al. 2015, 2022). This biased view may have important consequences for our overall understanding of cluster populations, since old-aged objects, including ancient globular clusters (GCs), are often more widely distributed in and around galaxies than young and intermediate-age ones (e.g., Rhode & Zepf 2004; Huxor et al. 2014; Veljanoski et al. 2015).
@@
While young to intermediate-age star clusters are of great interest for understanding star-formation processes and calibrating stellar evolution and population models (e.g., Grasha et al. 2019; Renzini & Fusi Pecci 1988; Dotter et al. 2007), old clusters provide probes of galaxy assembly histories (e.g., Brodie & Strader 2006). Evidence suggests many GCs are accreted through mergers, and their identification and analysis offer insights into the timeline and nature of hierarchical galaxy growth (e.g., Côté et al. 1998; Mackey et al. 2010). In the Milky Way (MW), searches for accreted GCs focus on cluster kinematics, chemistry and ages (e.g., Massari et al. 2019; Monty et al. 2024). In contrast, outside the MW, accreted GCs can often be recognised more directly through their association to tidal streams (e.g., Foster et al. 2014; Mackey et al. 2019). The GC populations of dwarf galaxies () are of particular interest not only for understanding how they might contribute to the build-up of more massive galaxy GC systems via accretion events, but also for understanding whether these small systems have acquired GCs from merger and accretion events themselves (Deason et al. 2014).
@@
In this paper, we use state-of-the-art data from the Euclid satellite (Euclid Collaboration: Mellier et al. 2025) to conduct an in-depth study of the star cluster populations in two Local Group (LG) dwarf irregular (dIrr) galaxies, NGC 6822 and IC 10. Aside from the Magellanic Clouds, these are the closest gas-rich dwarf galaxies to us and, with stellar masses of (1–4) (Pace 2024), they serve as valuable analogues to the star cluster-forming low-mass galaxies that are now being identified at high redshift (e.g., Mowla et al. 2024). Euclid’s spectacular capabilities allow us to conduct a comprehensive and systematic survey of star clusters in these systems that is unprecedented in its combination of depth, areal coverage, and spatial resolution. In particular, the pixel scale of Euclid corresponds to sub-parsec scales at the distances of these galaxies, while the FoV captures 6 kpc on a side. Additionally, Euclid extends star cluster studies in these systems to near-infrared (NIR) wavelengths. We summarise the properties of the two galaxies in Table 1 and briefly review their properties below.
@@ 2.1 NGC 6822
First identified by Barnard (1884), NGC 6822 is the third nearest dIrr galaxy after the Large and Small Magellanic Clouds and has a distance of 510 kpc determined from the tip of the red giant branch (RGB, Fusco et al. 2012). It has many luminous H ii regions (e.g., Hodge et al. 1988; Efremova et al. 2011) and studies suggest it has experienced an increase in its global star-formation rate (SFR) over the last few hundred Myr (Hodge 1980; Gallart et al. 1996; Fusco et al. 2014; Khatamsaz et al. 2024). NGC 6822’s young stellar component is predominantly concentrated within a central bar structure, aligned along the northsouth direction (Hodge 1977). In contrast, the older and intermediate-age populations exhibit elliptical distributions and extend to larger radii (e.g., Letarte et al. 2002; Battinelli et al. 2006; Tantalo et al. 2022). In addition, NGC 6822 contains a significant quantity of H i gas which lies in a disc that is misaligned to both the inner bar structure and the extended spheroid (de Blok & Walter 2000; Weldrake et al. 2003). Notably, this H i disc rotates perpendicular to the older stellar component and hence constitutes a dynamically decoupled structure (Demers et al. 2006). In spite of the complex stellar structure of NGC 6822, there is no compelling evidence for a recent merger or interaction (Zhang et al. 2021). In fact, most studies suggest that the galaxy is not even bound to the MW at the present epoch, although it may have entered within its virial radius in the past (e.g., Battaglia et al. 2022; Bennet et al. 2024).
@@
NGC 6822’s star cluster system has been well-studied, dating back to a century ago, although studies have been carried out in a somewhat piecemeal fashion. Using photographic plates, Hubble (1925) identified five potential clusters, and Hodge (1977) later confirmed four (Hubble-IV, VI, VII, and VIII) of these as genuine clusters, and identified 26 additional potential clusters in the main body. Follow-up studies of the Hubble (1925) clusters using HST and ground-based spectroscopy (Wyder et al. 2000; Chandar et al. 2000; Larsen et al. 2022; Hwang et al. 2014) confirmed Hubble-VII as a GC, Hubble-VI and VIII as likely young or intermediate-age clusters, and Hubble-IV to be associated with an H ii region. Krienke & Hodge (2004) further identified nine new cluster candidates in the central regions using HST imaging. In the remote halo, Hwang et al. (2011) identified four GCs using Canada–France–Hawaii telescope (CFHT) MegaCam data (), and Huxor et al. (2013) later discovered three more GCs from an independent analysis of the same data, plus additional coverage. Veljanoski et al. (2015) presented a homogeneous photometric and spectroscopic analysis of NGC 6288’s GC system and used it to infer a dynamical mass of (3–4) within 11 kpc, implying the system is very dark-matter dominated. Particularly striking is the linear alignment of the outermost GCs in NGC 6822, which extends over kpc at the distance of the galaxy.
@@ 2.2 IC 10
There is considerable uncertainty in the total reddening towards IC 10. Massey et al. (2007) measured the total reddening value to be using the location of the plume of blue supergiants from UBVRI wide-field imaging (). This aligns with the values of and reported by Sanna et al. (2008) and Richer et al. (2001), based on deep HST imaging of field stars and optical spectroscopy, respectively. It also agrees well with the average reddening value of from Demers et al. (2004), measured using Carbon stars over a field. However, their values vary across the field by as much as 0.30 magnitudes. A few higher values have also been reported, such as Kim et al. (2009), who found the total reddening to be through deep JHK photometry of the central field and more recently, using NIR and mid-IR data, Dell’Agli et al. (2018) found a value of , also restricted to a FoV that excludes the halo. The low Galactic latitude of IC 10 places it in the regime in which the foreground reddening maps of Schlegel et al. (1998) are known to be inaccurate. With the Schlafly & Finkbeiner (2011) recalibration, the dust maps return foreground values between 0.77 and 1.58, considerably larger than some of the total line-of-sight estimates in the literature.
Similarly to NGC 6822, IC 10’s metallicity has been estimated to be , based on its H ii regions (Skillman et al. 1989; Garnett 1990; Lee et al. 2003). IC 10’s older population has received less attention but using RGB stars, Tikhonov & Galazutdinova (2009) measured a photometric metallicity of .
Karachentsev & Tikhonov (1993) first identified seven star cluster candidates in IC 10 from ground-based B- and V-band images. Hunter (2001) later found additional candidates using HST imaging, focused on the central and eastern parts of the galaxy. Tikhonov & Galazutdinova (2009) conducted a more extensive search across the central body and parts of the halo using a combination of HST data and ground-based observations from the 6m BTA and 1m Zeiss telescopes, compiling a list of 57 candidates, subsuming those from Karachentsev & Tikhonov (1993) and some from Hunter (2001). Sharina et al. (2010) identified 17 new objects in archival HST data, while Lim & Lee (2015), hereafter L15, found five additional cluster candidates in the remote halo using Subaru R-band imaging with 07 seeing. As with NGC 6822, the previous cluster work has involved an amalgamation of patchy high-resolution HST coverage, wide-field low-resolution coverage, and varying search sensitivities across the galaxy.
@@ 3 Data
We combine the Euclid Early Release Observations (ERO, Cuillandre et al. 2025)111Euclid Early Release Observations (2024) with the NOAO Local Group Galaxy Survey (LGGS, Massey et al. 2007) to provide broad wavelength coverage from to , which we use for both star cluster identification and detailed characterisation of their properties.
@@ 4.1 Blind cluster search
The initial search was conducted by one of us (JH) on the images as these offer the highest spatial resolution compared to the NIR filters, and a greater sensitivity to various cluster ages due to the broad filter bandwidth. The extended-emission images were preferred as they help to highlight instances where there is significant coincident nebular emission. We excluded cases where an object is so heavily embedded in nebular emission that it becomes difficult to discern whether it is a distinct star cluster or a sparser OB association. The search was conducted blindly, without prior knowledge of existing cluster locations, though we inspected cutouts of a few already confirmed clusters before starting, to familiarise ourselves with the likely range of morphologies. This was helpful for learning how to identify faint clusters and for calibrating our expectations for the appearance of star clusters at these distances, and with Euclid resolution.
Once the cluster candidates were identified in , they were inspected across the other eight bands: U, B, V, R, and I, as well as , , . This aided in distinguishing between clusters and background galaxies, as the galaxies tend to drop out in the shorter-wavelength passbands. Objects exhibiting this behaviour, or showing other indications of a not being a star cluster, were immediately discarded. Remaining candidates were labelled as either ?likely? or ?maybe? and were passed for further classification by two additional team members (AF and SL). By the end of this process, each object initially classed as ?maybe? or ?likely? also had two further (independent) classification attempts into the categories of ?likely?,?maybe?, and ?unlikely?. The final classification of a given star cluster candidate was arrived at as follows:
Three ?likely? classifications == 1;
Two ?likely? and one ?maybe? == 2;
One ?likely? and two ?maybe? == 3;
Three ?maybe? == 4;
Any combination with ?unlikely? == 5.
@@ 4.2 Literature sample compilation
As mentioned before, previous studies have searched for star clusters in NGC 6822 and IC 10 and we sought to identify which of the candidates uncovered in our Euclid blind search were already known. To this end, we undertook a thorough literature source compilation. The process was mostly straightforward for IC 10, due to the list of 66 cluster candidates already compiled from various sources (Karachentsev & Tikhonov 1993; Hunter 2001; Sharina et al. 2010) by L15. However, we discovered that 21 previously identified cluster candidates were missing from the list by L15 (four from Hunter 2001 and 17 from Sharina et al. 2010), which we included in our compilation.
@@
In the case of NGC 6822, we did not have the benefit of a recent compilation and hence had to tabulate directly the cluster candidates presented in multiple works spanning several decades (Hubble 1925; Hodge 1977; Krienke & Hodge 2004; Hwang et al. 2011; Huxor et al. 2013).444We note that Gouliermis et al. (2010) studied hierarchical clustering of blue stars in NGC 6822 using ground-based imaging. Using a nearest-neighbour density method, they defined clusters as overdensities of point sources that are above the average stellar density. However, apart from C24 from Hodge (1977), these clusters are not visually apparent in the Euclid image. Their clusters span a large range of sizes (10–120 pc), with most of them having half-light radii pc, which is considerably larger than the objects we identify as star cluster candidates in this paper. Indeed, combined with their stellar masses of –, these objects are more akin to low-mass unbound stellar associations. We therefore exclude the Gouliermis et al. (2010) candidates from our final lists and do not aim for completeness in very low-mass young clusters. In older studies, cluster candidates generally lacked published coordinates (Hodge 1977) or had coordinates that did not always correspond to a visible overdensity in the Euclid images (Krienke & Hodge 2004). In such cases, positions were determined on a best-effort basis, either by updating the coordinates to match the nearest cluster candidate to the provided coordinates or, for Hodge (1977), the closest cluster candidate to the position highlighted in their finding chart.
@@
In total, we compiled a list of 87 previously identified cluster candidates within IC 10 and 43 within NGC 6822 (noting an additional three NGC 6822 clusters outside of the Euclid footprint). Cross-matching the list of candidates from our blind search with those from the literature, we found that we recovered many but not all of the sources identified in earlier studies, as well as adding many new sources not previously known. For completeness, all three inspectors examined the positions of the literature candidates and classified them using the same system described above. Reassuringly, all the literature candidates within the IC 10 and NGC 6822 footprints that we classified as classes 1 and 2 already appeared in our blind search list. On the other hand, several literature candidates were missed. In some cases, the reported positions did not correspond to any obvious stellar overdensity in the Euclid images. In other cases, the reported positions clearly corresponded to single stars, nebulae, or background galaxies; such objects were assigned to class 6, indicating sources that are definitely not star clusters. In NGC 6822, our blind search missed 23 out of 43 literature candidates. Of these, one was classified as class 3 and one as class 4, while 14 were assigned to class 5 and four to class 6. As a result, the majority were deemed unconvincing as cluster candidates. In IC 10, we missed 30 out of 87 previous candidates. Of these, one was classified as class 3, six as class 4, 18 as class 5, and five as class 6. Notably, of the five candidate clusters in the remote halo of IC 10 found by L15 using Subaru R-band imaging, three were recovered in our blind search and are confirmed as genuine clusters (ESCC-IC10-49, 51 and 52; 49 and 51 are displayed in Fig. 2), while the remaining two (L64 and L65) are clearly background galaxies and have been assigned to class 6. While it is unlikely that many class 5 sources are genuine clusters, we note that objects previously identified as faint clusters at UV wavelengths may appear less convincing in the filter due to its longer wavelength. Similarly, it is possible that some faint sources previously identified at HST resolution may be less obvious at the slightly lower resolution of Euclid. However, since class 5 sources do not appear cluster-like in the Euclid data, we do not analyse them further in this work. For completeness and to facilitate future studies of these cluster systems, we include all literature candidates, whether recovered or not, and regardless of our classification, in Tables 2 and 3, retaining their original identifiers (based on the earliest published ID). These objects can be identified as having an entry in the column ?Recovered??. Thumbnails of the previously-identified candidates that we reclassify as class 5 and 6 are included in the Appendix (Figs. 20 and 21), alongside all other previously identified and newly discovered clusters.
@@
In summary, our blind search identified 30 new star cluster candidates of classes 1–4 in NGC 6822 and 16 in IC 10. Incorporating the 22 and 55 previously reported literature candidates with classifications 1–4 (of which 20 and 48 were independently recovered in our blind search), the total number of candidates in these classes increases to 52 in NGC 6822 and 71 in IC 10. For clarity, we refer to our final list of class 1–4 cluster candidates simply as ?clusters? throughout the remainder of the paper.

### v3

@@
Star clusters are valuable indicators of galaxy evolution, offering insights into the buildup of stellar populations across cosmic time. Understanding the intrinsic star cluster populations of dwarf galaxies is particularly important given these systems’ role in the hierarchical growth of larger systems. We use data from Euclid’s Early Release Observation programme to study star clusters in two star-forming dwarf irregular galaxies in the Local Group, NGC 6822 and IC 10 [ (1–4) ]. With Euclid, star clusters are resolved into individual stars across the main bodies and haloes of both galaxies. Through visual inspection of the images, we uncover 30 new star cluster candidates in NGC 6822 and 16 in IC 10, ranging from compact to diffuse extended clusters. We compile and re-evaluate previously identified literature candidates, resulting in final combined catalogues of 52 (NGC 6822) and 71 (IC 10) cluster candidates with confidence-based classifications. We present homogeneous photometry in , , , and , and in archival UBVRI data, alongside size measurements and properties derived from the spectral energy distribution fitting code BAGPIPES. Through synthetic cluster injection, we conclude our sample is complete to for ages , and to for ages of . We find that IC 10 has more young clusters than NGC 6822, and its young clusters extend to higher masses, consistent with its starburst nature. We find several old massive () clusters in both dwarfs, including an exceptional cluster in NGC 6822’s outskirts with a mass of , nearly twice as massive as any other old cluster in either galaxy. In NGC 6822, we also identify a previously undetected, old, and extended cluster (). Using well-defined criteria, we identify 11 candidate GCs in NGC 6822 and nine in IC 10. Both galaxies have high specific frequencies () for their luminosities but remain consistent with the known GC scaling relationships in the low-luminosity regime.
@@
Star clusters are present in nearly all galaxies and can be observed across most of cosmic time (e.g. Forbes et al. 2018a; Adamo et al. 2024). They are likely the sites of most, if not all, star formation in the Universe, even if they do not always survive for long as coherent or bound groupings of stars. In the Local Volume, star clusters are observed with a remarkable range of properties, from massive and young clusters (e.g. Portegies Zwart et al. 2010) to ancient, faint, and extended ones (e.g. Huxor et al. 2005; Crnojević et al. 2016). The origin of this diversity is not fully understood but likely reflects the complex interplay between birth conditions, stellar feedback, and dynamical evolution (e.g. Krumholz et al. 2019; Lahén et al. 2025). Nearby galaxies offer key advantages for studying star clusters, as their proximity allows resolution into individual stars, facilitating both identification and characterisation (e.g. Huxor et al. 2014; Johnson et al. 2017; Jang et al. 2012). Furthermore, using resolved stars, star clusters can be observed to lower luminosities and surface densities than is possible in more distant systems, yielding insight into small-scale star-formation events and the bottom end of the cluster mass function.
@@
Comprehensive views of star cluster populations in nearby galaxies are still relatively rare. Historically, many extragalactic star cluster studies have been conducted with the Hubble Space Telescope (HST), which captures only a small field of view (FoV) in a single pointing. Studies often choose to focus on the bright inner regions of galaxies where recent star formation is concentrated and hence the number of young clusters is highest (e.g. Adamo et al. 2017). Even large HST imaging campaigns, such as the PHAT and PHATTER surveys of M31 and M33, capture star clusters only in the main disc (Johnson et al. 2015, 2022). This biased view may have important consequences for our overall understanding of cluster populations, since old-aged objects, including ancient globular clusters (GCs), are often more widely distributed in and around galaxies than young and intermediate-age ones (e.g. Rhode and Zepf 2004; Huxor et al. 2014; Veljanoski et al. 2015).
@@
While young to intermediate-age star clusters are of great interest for understanding star-formation processes and calibrating stellar evolution and population models (e.g. Grasha et al. 2019; Renzini and Fusi Pecci 1988; Dotter et al. 2007), old clusters provide probes of galaxy assembly histories (e.g. Brodie and Strader 2006). Evidence suggests many GCs are accreted through mergers, and their identification and analysis offer insights into the timeline and nature of hierarchical galaxy growth (e.g. Côté et al. 1998; Mackey et al. 2010). In the Milky Way (MW), searches for accreted GCs focus on cluster kinematics, chemistry and ages (e.g. Massari et al. 2019; Monty et al. 2024). In contrast, outside the MW, accreted GCs can often be recognised more directly through their association to tidal streams (e.g. Foster et al. 2014; Mackey et al. 2019). The GC populations of dwarf galaxies () are of particular interest not only for understanding how they might contribute to the build-up of more massive galaxy GC systems via accretion events, but also for understanding whether these small systems have acquired GCs from merger and accretion events themselves (Deason et al. 2014).
@@
In this paper, we use state-of-the-art data from the Euclid satellite (Euclid Collaboration: Mellier et al. 2025) to conduct an in-depth study of the star cluster populations in two Local Group (LG) dwarf irregular (dIrr) galaxies, NGC 6822 and IC 10. Aside from the Magellanic Clouds, these are the closest gas-rich dwarf galaxies to us and, with stellar masses of (1–4) (Pace 2024), they serve as valuable analogues to the star cluster-forming low-mass galaxies that are now being identified at high redshift (e.g. Mowla et al. 2024). Euclid’s spectacular capabilities allow us to conduct a comprehensive and systematic survey of star clusters in these systems that is unprecedented in its combination of depth, areal coverage, and spatial resolution. In particular, the pixel scale of Euclid corresponds to sub-parsec scales at the distances of these galaxies, while the FoV captures 6 kpc on a side. Additionally, Euclid extends star cluster studies in these systems to near-infrared (NIR) wavelengths. We summarise the properties of the two galaxies in Table 1 and briefly review their properties below.
@@ 2.1 NGC 6822
First identified by Barnard (1884), NGC 6822 is the third nearest dIrr galaxy after the Large and Small Magellanic Clouds and has a distance of 510 kpc determined from the tip of the red giant branch (RGB, Fusco et al. 2012). It has many luminous H ii regions (e.g. Hodge et al. 1988; Efremova et al. 2011) and studies suggest it has experienced an increase in its global star-formation rate (SFR) over the last few hundred Myr (Hodge 1980; Gallart et al. 1996; Fusco et al. 2014; Khatamsaz et al. 2024). NGC 6822’s young stellar component is predominantly concentrated within a central bar structure, aligned along the northsouth direction (Hodge 1977). In contrast, the older and intermediate-age populations exhibit elliptical distributions and extend to larger radii (e.g. Letarte et al. 2002; Battinelli et al. 2006; Tantalo et al. 2022). In addition, NGC 6822 contains a significant quantity of H i gas which lies in a disc that is misaligned to both the inner bar structure and the extended spheroid (de Blok and Walter 2000; Weldrake et al. 2003). Notably, this H i disc rotates perpendicular to the older stellar component and hence constitutes a dynamically decoupled structure (Demers et al. 2006). In spite of the complex stellar structure of NGC 6822, there is no compelling evidence for a recent merger or interaction (Zhang et al. 2021). In fact, most studies suggest that the galaxy is not even bound to the MW at the present epoch, although it may have entered within its virial radius in the past (e.g. Battaglia et al. 2022; Bennet et al. 2024).
@@
NGC 6822’s star cluster system has been well-studied, dating back to a century ago, although studies have been carried out in a somewhat piecemeal fashion. Using photographic plates, Hubble (1925) identified five potential clusters, and Hodge (1977) later confirmed four (Hubble-IV, VI, VII, and VIII) of these as genuine clusters, and identified 26 additional potential clusters in the main body. Follow-up studies of the Hubble (1925) clusters using HST and ground-based spectroscopy (Wyder et al. 2000; Chandar et al. 2000; Larsen et al. 2022; Hwang et al. 2014) confirmed Hubble-VII as a GC, Hubble-VI and VIII as likely young or intermediate-age clusters, and Hubble-IV to be associated with an H ii region. Krienke and Hodge (2004) further identified nine new cluster candidates in the central regions using HST imaging. In the remote halo, Hwang et al. (2011) identified four GCs using Canada–France–Hawaii telescope (CFHT) MegaCam data (), and Huxor et al. (2013) later discovered three more GCs from an independent analysis of the same data, plus additional coverage. Veljanoski et al. (2015) presented a homogeneous photometric and spectroscopic analysis of NGC 6288’s GC system and used it to infer a dynamical mass of (3–4) within 11 kpc, implying the system is very dark-matter dominated. Particularly striking is the linear alignment of the outermost GCs in NGC 6822, which extends over kpc at the distance of the galaxy.
@@ 2.2 IC 10
There is considerable uncertainty in the total reddening towards IC 10. Massey et al. (2007) measured the total reddening value to be using the location of the plume of blue supergiants from UBVRI wide-field imaging (). This aligns with the values of and reported by Sanna et al. (2008) and Richer et al. (2001), based on deep HST imaging of field stars and optical spectroscopy, respectively. It also agrees well with the average reddening value of from Demers et al. (2004), measured using Carbon stars over a field. However, their values vary across the field by as much as 0.30 magnitudes. A few higher values have also been reported, such as Kim et al. (2009), who found the total reddening to be through deep JHK photometry of the central field and more recently, using NIR and mid-IR data, Dell’Agli et al. (2018) found a value of , also restricted to a FoV that excludes the halo. The low Galactic latitude of IC 10 places it in the regime in which the foreground reddening maps of Schlegel et al. (1998) are known to be inaccurate. With the Schlafly and Finkbeiner (2011) recalibration, the dust maps return foreground values between 0.77 and 1.58, considerably larger than some of the total line-of-sight estimates in the literature.
Similarly to NGC 6822, IC 10’s metallicity has been estimated to be , based on its H ii regions (Skillman et al. 1989; Garnett 1990; Lee et al. 2003). IC 10’s older population has received less attention but using RGB stars, Tikhonov and Galazutdinova (2009) measured a photometric metallicity of .
Karachentsev and Tikhonov (1993) first identified seven star cluster candidates in IC 10 from ground-based B- and V-band images. Hunter (2001) later found additional candidates using HST imaging, focused on the central and eastern parts of the galaxy. Tikhonov and Galazutdinova (2009) conducted a more extensive search across the central body and parts of the halo using a combination of HST data and ground-based observations from the 6m BTA and 1m Zeiss telescopes, compiling a list of 57 candidates, subsuming those from Karachentsev and Tikhonov (1993) and some from Hunter (2001). Sharina et al. (2010) identified 17 new objects in archival HST data, while Lim and Lee (2015), hereafter L15, found five additional cluster candidates in the remote halo using Subaru R-band imaging with 07 seeing. As with NGC 6822, the previous cluster work has involved an amalgamation of patchy high-resolution HST coverage, wide-field low-resolution coverage, and varying search sensitivities across the galaxy.
@@ 3 Data
We combined the Euclid Early Release Observations (ERO, Cuillandre et al. 2025)111Euclid Early Release Observations (2024) with the NOAO Local Group Galaxy Survey (LGGS, Massey et al. 2007) to provide broad wavelength coverage from to , which we used for both star cluster identification and detailed characterisation of their properties.
@@ 4.1 Blind cluster search
The initial search was conducted by one of us (JH) on the images as these offer the highest spatial resolution compared to the NIR filters, and a greater sensitivity to various cluster ages due to the broad filter bandwidth. The extended-emission images were preferred as they help to highlight instances where there is significant coincident nebular emission. We excluded cases where an object is so heavily embedded in nebular emission that it becomes difficult to discern whether it is a distinct star cluster or a sparser OB association. We also note that very red clusters (heavily dust reddened or embedded) may be more clearly distinguished in the NIR and therefore underrepresented by our search. The search was conducted blindly, without prior knowledge of existing cluster locations, though we inspected cutouts of a few already confirmed clusters before starting, to familiarise ourselves with the likely range of morphologies. This was helpful for learning how to identify faint clusters and for calibrating our expectations for the appearance of star clusters at these distances, and with Euclid resolution. Once the cluster candidates were identified in , they were inspected across the other eight bands: U, B, V, R, and I, as well as , , . This aided in distinguishing between clusters and background galaxies, as the galaxies tend to drop out in the shorter-wavelength passbands. Objects exhibiting this behaviour, or showing other indications of a not being a star cluster, were immediately discarded. Remaining candidates were labelled as either ‘likely’ or ‘maybe’ and were passed for further classification by two additional team members (AF and SL). By the end of this process, each object initially classed as maybe or likely also had two further (independent) classification attempts into the categories of likely, maybe, and ‘unlikely’. The final classification of a given star cluster candidate was arrived at as follows:
Three likely classifications == 1;
Two likely and one maybe == 2;
One likely and two maybe == 3;
Three maybe == 4;
Any combination with unlikely == 5.
@@ 4.2 Literature sample compilation
As mentioned before, previous studies have searched for star clusters in NGC 6822 and IC 10 and we sought to identify which of the candidates uncovered in our Euclid blind search were already known. To this end, we undertook a thorough literature source compilation. The process was mostly straightforward for IC 10, due to the list of 66 cluster candidates already compiled from various sources (Karachentsev and Tikhonov 1993; Hunter 2001; Sharina et al. 2010) by L15. However, we discovered that 21 previously identified cluster candidates were missing from the list by L15 (four from Hunter 2001 and 17 from Sharina et al. 2010), which we included in our compilation.
@@
In the case of NGC 6822, we did not have the benefit of a recent compilation and hence had to tabulate directly the cluster candidates presented in multiple works spanning several decades (Hubble 1925; Hodge 1977; Krienke and Hodge 2004; Hwang et al. 2011; Huxor et al. 2013).444We note that Gouliermis et al. (2010) studied hierarchical clustering of blue stars in NGC 6822 using ground-based imaging. Using a nearest-neighbour density method, they defined clusters as overdensities of point sources that are above the average stellar density. However, apart from C24 from Hodge (1977), these clusters are not visually apparent in the Euclid image. Their clusters span a large range of sizes (10–120 pc), with most of them having half-light radii pc, which is considerably larger than the objects we identify as star cluster candidates in this paper. Indeed, combined with their stellar masses of –, these objects are more akin to low-mass unbound stellar associations. We therefore exclude the Gouliermis et al. (2010) candidates from our final lists and do not aim for completeness in very low-mass young clusters. In older studies, cluster candidates generally lacked published coordinates (Hodge 1977) or had coordinates that did not always correspond to a visible overdensity in the Euclid images (Krienke and Hodge 2004). In such cases, positions were determined on a best-effort basis, either by updating the coordinates to match the nearest cluster candidate to the provided coordinates or, for Hodge (1977), the closest cluster candidate to the position highlighted in their finding chart.
@@
In total, we compiled a list of 87 previously identified cluster candidates within IC 10 and 43 within NGC 6822 (noting an additional three NGC 6822 clusters outside of the Euclid footprint). Cross-matching the list of candidates from our blind search with those from the literature, we found that we recovered many but not all of the sources identified in earlier studies, as well as adding many new sources not previously known. For completeness, all three inspectors examined the positions of the literature candidates and classified them using the same system described above. Reassuringly, all the literature candidates within the IC 10 and NGC 6822 footprints that we classified as classes 1 and 2 already appeared in our blind search list. On the other hand, several literature candidates were missed. In some cases, the reported positions did not correspond to any obvious stellar overdensity in the Euclid images. In other cases, the reported positions clearly corresponded to single stars, nebulae, or background galaxies; such objects were assigned to class 6, indicating sources that are definitely not star clusters. In NGC 6822, our blind search missed 23 out of 43 literature candidates. Of these, one was classified as class 3 and one as class 4, while 14 were assigned to class 5 and four to class 6. As a result, the majority were deemed unconvincing as cluster candidates. In IC 10, we missed 30 out of 87 previous candidates. Of these, one was classified as class 3, six as class 4, 18 as class 5, and five as class 6. Notably, of the five candidate clusters in the remote halo of IC 10 found by L15 using Subaru R-band imaging, three were recovered in our blind search and are confirmed as genuine clusters (ESCC-IC10-49, 51 and 52; 49 and 51 are displayed in Fig. 2), while the remaining two (L64 and L65) are clearly background galaxies and have been assigned to class 6. While it is unlikely that many class 5 sources are genuine clusters, we note that objects previously identified as faint clusters at UV wavelengths may appear less convincing in the filter due to its longer wavelength. Similarly, it is possible that some faint sources previously identified at HST resolution may be less obvious at the slightly lower resolution of Euclid. However, since class 5 sources do not appear cluster-like in the Euclid data, we do not analyse them further in this work. For completeness and to facilitate future studies of these cluster systems, we include all literature candidates, whether recovered or not, and regardless of our classification, in Tables 2 and 3, retaining their original identifiers (based on the earliest published ID). These objects can be identified as having an entry in the column ‘Recovered?’. Thumbnails of the previously-identified candidates that we reclassify as class 5 and 6 are included in the Appendix (Figs. 20 and 21), alongside all other previously identified and newly discovered clusters.
@@
In summary, our blind search identified 30 new star cluster candidates of classes 1–4 in NGC 6822 and 16 in IC 10. Incorporating the 22 and 55 previously reported literature candidates with classifications 1–4 (of which 20 and 48 were independently recovered in our blind search), the total number of candidates in these classes increases to 52 in NGC 6822 and 71 in IC 10. For clarity, we refer to our final list of class 1–4 cluster candidates simply as ‘clusters’ throughout the remainder of the paper.